OpenStudy (anonymous):

PLEASE HELP!! The only one I don't understand. Decide which part of the quadratic formula tells you whether the quadratic equation can be solved by factoring. −b b^2 − 4ac 2a Use the part of the quadratic formula that you chose above and find its value given the following quadratic equation: 2x2 + 7x + 3 = 0 Numerical Answers Expected!

OpenStudy (anonymous):

sqrt(b^2-4ac) part

OpenStudy (anonymous):

you have 2x^2+7x+3=0 which is similar to ax^2+bx+c=0 So what's b^2-4ac=?